6. Assume that an electron is located somewhere within a region of atomic size. Estimate the minimum uncertainty in its momentum. By assuming that this uncertainty is comparable with its average momentum, estimate the average kinetic energy of the electron.
1
Expert's answer
2020-06-18T10:59:17-0400

According to the Heisenberg's uncertainty principle, the minimum uncertainty in momentum will be:


"\Delta p = \dfrac{\hbar}{2\Delta x}"

where "\hbar = 1.05\cdot 10^{-34}J\cdot s" is the reduced Planck constant and "\Delta x = 1.06\cdot 10^{-10}m" is the atom diameter (uncertainty of the coordinate). Thus:


"\Delta p = \dfrac{1.05\cdot 10^{-34}J\cdot s}{2\cdot 1.06\cdot 10^{-10}m} = 5\cdot 10^{-25} kg\cdot m/s"

The average kinetic energy of the electron will be:


"K = \dfrac{\Delta p^2}{2m_e}"

where "m_e = 9.11\cdot 10^{-31}kg" is the mass of electon.


"K = \dfrac{(5^\cdot 10^{-25} kg\cdot m/s)^2}{2\cdot 9.11\cdot 10^{-31}kg} = 1.37\cdot 10^{19}J"